total pressure of a gas mixture composed of 0.622 atm He gas, 844 torr Ne gas and 378 mmHg Ar gas?
Answer #1

We know the Dalton's law

PTotal  = P1 + P2 + P3 + ...............Pn

Total pressure is equal to sum of partial pressure

write down all gas pressure in same unit

PHe = 0.622 atm = 472.7 mmHg (1 atm = 760 mmHg then 0.622 atm = 0.622 X 760 = 472.7 mmHg)

PNe = 344 torr = 344 mmHg (1 torr = 1 mmHg then 344 torr = 344 mmHg )

PAr = 378 mmHg

PTotal  = 472.7 mmHg + 344 mmHg + 378 mHg = 1194.7 mm Hg

Total pressure = 1194.7 mmHg

or

Total pressure = 1.572 atm (1 atm = 760 mmHg then 1194.7 mmHg  = 1194.7 / 760 = 1.572 atm)

or

Total pressure = = 1194.7 torr (1 torr = 1 mmHg then 1194.7 mmHg  = 1194.7 torr )